OpenStudy (anonymous):

Point E is located at (2, −3), and point F is located at (−2, −1). Find the y value for the point that is 3 over 4 the distance from point E to point F. 1.75 0.5 −2 −1.5

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

2. Find the y value for the point that splits segment EF in half if point E is located at (5, −2) and point F is located at (−1, 3). 1 0.5 4.5 2

OpenStudy (anonymous):

3. Line AB contains points A (−2, 3) and B (4, 5). Line AB has a slope that is zero undefined positive negative

OpenStudy (bradely):

y= (my2+ny1)/(m+n) x= (3(-1)+4(-3))/(3+4) =(-3-12)/7 =-15/7 =-2.1 approx =-2 option 3 is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

i got 1.71 @bradely

OpenStudy (anonymous):

3/4 becomes 3/7 because its a ratio...and 2- (-2) because their x's... so i multiplied 3/7*4/1 =1.71